OpenStudy (anonymous):

Use the definition of continuity and the properties of limits to show that the function is continuous at the given number a. f(x)= (x+2x^3)^4 , a=-1

OpenStudy (anonymous):

can someone help me please!

OpenStudy (anonymous):

The function will be continuous at \(a\) if \[\lim_{x\to a}f(x)=f(a)\] We have that \[f(a)=f(-1)=\left(-1+2(-1)^3\right)^4=(-1-2)^4=81\] If the limit is the same, the function is continuous.

OpenStudy (anonymous):

:) thank you so much

OpenStudy (anonymous):

yw